# Flexible Physical Camouflage Generation Based on a Differential Approach

### v1

@@ 4.6 Ablation Studies
In this section of the experiment, we investigated the effects of different training methods (including diffusion model modules, and initialization areas) and loss function terms on adversarial camouflage.
@@
Different Combinations of Loss Functions. We investigated the impact of different combinations of loss functions on adversarial camouflage, including ${L_{attack}}$, ${L_{attack}+L_{vt}}$, and ${L_{attack}+L_{nps}}$ for validation. Meanwhile, we compared the AP@0.5 computed on the test set collected based on Carla. The experimental results are presented in Table.8, indicating a decrease in the performance of adversarial camouflage in the absence of certain loss function terms. When considering only the ${L_{attack}}$ term, the attack effectiveness surpasses the other two experimental groups. However, as depicted in Figure.13, the absence of the ${L_{vt}}$ term leads to numerous black spots in the camouflage, resulting in reduced concealment. Conversely, the lack of the ${L_{nps}}$ constraint results in brighter camouflage colors, leading to significant pixel information loss during printing. Therefore, by combining ${L_{attack}}$, ${L_{nps}}$, and ${L_{vt}}$, we achieve stronger attack performance and better concealment for the adversarial camouflage.
@@
Table 8: The comparison results of AP@0.5 under different loss function combination schemes.
Method
YOLOv3 YOLOv5
FRCNN
DETR
${L_{attack}}$
0.286 0.226
0.442
0.284
${L_{attack}}+L_{vt}$
0.626 0.534
0.612
0.780
${L_{attack}}+L_{nps}$
0.506 0.281
0.468
0.416
${L_{attack}}+L_{vt}+L_{nps}$
0.159 0.088
0.319
0.286
@@
Different Training Methods In this section, we investigate the impact of the diffusion model on optimizing against camouflage and the effect of the initialization area on the attack’s effectiveness. We establish two optimization methods: traditional gradient-based optimization and diffusion model-based optimization as described in Section 3.2. Meanwhile, two methods of initial area change are set, local coverage and global coverage. Local coverage refers to partial regions on structures such as car doors and roofs, while global coverage refers to the camouflage covering the entire vehicle (excluding structures such as windows and tires). Specifically, based on the training data collected from Carla and using YOLOv5 as the victim detection model, we generated four sets of adversarial camouflages:
@@
Where Diffuse-Global represents the FPA(v5) adversarial camouflage used in the previous comparison. The AP@0.5 results of the adversarial camouflage under different detection models are presented in Table.9. It is observed that the diffusion model optimization approach significantly outperforms the gradient approach. Meanwhile, enlarging the coverage area significantly enhances the effectiveness of camouflage. As observed in Figure.13, both gradient and diffusion model optimization-based camouflages effectively conceal the car, underscoring the robustness of our training methodology and the design of loss functions.
Table 9: Comparison of the AP@0.5 results of adversarial camouflages generated by different optimization methods and initial areas under different detection models.
Method YOLOv3 YOLOv5
FRCNN
DETR
Gradient-Local 0.676 0.565
0.681
0.785
Gradient-Global 0.527 0.493
0.605
0.664
Diffuse-Local 0.481 0.415
0.536
0.436
Diffuse-Global 0.159 0.088
0.319
0.286
5 Conclusion
In conclusion, our work presents a novel approach to adversarial camouflage through the integration of a comprehensive 3D rendering framework, departing from conventional neural rendering methods. Leveraging the inherent capabilities of the framework, we achieved a nuanced representation of 3D targets by simulating authentic lighting conditions and material variations, thereby contributing to a more realistic rendering outcome. The combination of a generatively learned adversarial pattern from diffusion models, along with a designed adversarial loss and covert constraint loss, ensures the adversarial and covert nature of the camouflage in the physical world. The contribution of our work lies in the establishment of a rendering system rooted in a robust 3D framework, which enhances rendering efficiency and enables the generation of adversarial camouflage in sticker mode for practical deployment. The incorporation of confrontation loss and concealment constraint loss further strengthens the aggressiveness and concealment of the camouflage, demonstrating adaptability across diverse environments. In future works, we aim to explore additional refinements to the adversarial patterns, investigate real-world deployment scenarios, and evaluate the generalization of our approach to different objects and environments, thereby advancing the field of adversarial camouflage in both theoretical and practical dimensions.
Acknowledgments
This research is supported by the xxx.

### v3

@@ IV-F Ablation Studies
In this section of the experiment, we investigated the effects of different training methods (including diffusion model modules and initialization areas) and terms of loss function on adversarial camouflage.
@@
Different Combinations of Loss Functions. We investigated the impact of different combinations of loss functions on adversarial camouflage, including ${L_{attack}}$, ${L_{attack}+L_{vt}}$, and ${L_{attack}+L_{nps}}$ for validation. Meanwhile, we compared the AP@0.5 computed on the test set collected based on Carla. The experimental results are presented in Table.VIII, indicating a decrease in the performance of adversarial camouflage in the absence of certain terms of the loss function. When considering only the ${L_{attack}}$ term, the attack effectiveness surpasses the other two experimental groups. However, as depicted in Figure.14, the absence of the ${L_{vt}}$ term leads to numerous black spots in the camouflage, resulting in reduced concealment. In contrast, the lack of the constraint ${L_{nps}}$ results in brighter camouflage colors, leading to significant loss of information of the pixels during printing. Therefore, combining ${L_{attack}}$, ${L_{nps}}$, and ${L_{vt}}$, we achieve stronger attack performance and better concealment for adversarial camouflage.
@@
TABLE VIII: The comparison results of AP@0.5 under different loss function combination schemes.
Method YOLOv3 YOLOv5 FRCNN DETR
${L_{attack}}$ 0.286 0.226 0.442 0.284
${L_{attack}}+L_{vt}$ 0.626 0.534 0.612 0.780
${L_{attack}}+L_{nps}$ 0.506 0.281 0.468 0.416
${L_{attack}}+L_{vt}+L_{nps}$ 0.159 0.088 0.319 0.286
@@
Different Training Methods In this section, we investigate the impact of the diffusion model on optimizing against camouflage and the effect of the initialization area on the attack’s effectiveness. We establish two optimization methods: traditional gradient-based optimization and diffusion model-based optimization. Meanwhile, two methods of initial area change are set, local coverage and global coverage. Local coverage refers to partial regions on structures such as car doors and roofs, while global coverage refers to the camouflage covering the entire vehicle (excluding structures such as windows and tires). Specifically, based on the training data collected from Carla and using YOLOv5 as the victim detection model, we generated four sets of adversarial camouflages:
@@
Where Diffuse-Global represents the FPA(v5) adversarial camouflage used in the previous comparison. The AP@0.5 results of the adversarial camouflage under different detection models are presented in Table.IX. It is observed that the diffusion model optimization approach significantly outperforms the gradient approach. Meanwhile, enlarging the coverage area significantly enhances the effectiveness of camouflage. As observed in Figure.14, both gradient and diffusion model optimization-based camouflages effectively conceal the car, underscoring the robustness of our training methodology and the design of loss functions.
TABLE IX: Comparison of the AP@0.5 results of adversarial camouflages generated by different optimization methods and initial areas under different detection models.
Method YOLOv3 YOLOv5 FRCNN DETR
Gradient-Local 0.676 0.565 0.681 0.785
Gradient-Global 0.527 0.493 0.605 0.664
Diffuse-Local 0.481 0.415 0.536 0.436
Diffuse-Global 0.159 0.088 0.319 0.286
V Discussions
Implications In conclusion, our work presents a novel approach to adversarial camouflage through the integration of a comprehensive 3D rendering framework, departing from conventional neural rendering methods. Using the inherent capabilities of the framework, we achieved a nuanced representation of 3D targets by simulating authentic lighting conditions and material variations, thereby contributing to a more realistic rendering outcome. The combination of a generatively learned adversarial pattern from diffusion models, along with a designed adversarial loss and covert constraint loss, ensures the adversarial and covert nature of the camouflage in the physical world. The contribution of our work lies in the establishment of a rendering system rooted in a robust 3D framework, which enhances rendering efficiency and enables the generation of adversarial camouflage in sticker mode for practical deployment. The incorporation of confrontation loss and concealment constraint loss further strengthens the aggressiveness and concealment of the camouflage, demonstrating adaptability across diverse environments. In future works, we aim to explore additional refinements to adversarial patterns, investigate real-world deployment scenarios, and evaluate the generalization of our approach to different objects and environments, thereby advancing the field of adversarial camouflage in both theoretical and practical dimensions.
Limitations Due to the limitations of the renderer’s capabilities, FPA is unable to simulate deformation during the rendering process. Consequently, when dealing with non-rigid targets, the effectiveness of the generated adversarial camouflage in real-world physical tests may exhibit significant deviations. In the future, we will further investigate adversarial texture generation techniques for 3D non-rigid targets.
Acknowledgment
This work is funded by the National Natural Science Foundation of China (No.62103330, 62233014).